OpenStudy (anonymous):

dia was a famous greek mathematician who lived and worked in alexandria egypt. He was a boy for 1/6 of his life. after 1/12 more, he acquired a beard. After 1/7 he married,in the fifth year of marriage his son was born. the son lived half as many years as dia . dia died four years after his son. how old was dia when he died?

OpenStudy (anonymous):

Hell of a question.

OpenStudy (anonymous):

yes it is lol

OpenStudy (anonymous):

If i remember correctly, this was actually put on the tombstone of a mathematician :)

OpenStudy (anonymous):

ask dia when she died

OpenStudy (akshay_budhkar):

OMG!!! what a question! my pen's hot trying it

OpenStudy (anonymous):

i know im still working it

OpenStudy (anonymous):

good to know its a challenge for others as well

OpenStudy (anonymous):

wait man i am working

OpenStudy (lalaly):

Final age = child + youth + bach + marriage + son's age + 4 x =1/6 x +1/12 x +1/7 x + 5 + 1/2x + 4 x=75/84 x +9 x=84

OpenStudy (anonymous):

Oh damn.

OpenStudy (anonymous):

COULD IT BE?

OpenStudy (anonymous):

um i didnt get that lalaly

OpenStudy (anonymous):

yeah one problem

OpenStudy (dumbcow):

Let d be age of dia, s be age of dia's son d = 2s son was born in 5th year of marriage which is 1/6+1/7+1/12 = 99/252 = into dia life son dies 4 years before dia s = (d-4) - (99/252d +5) s = 29/42d -9 By substitution d/2 = 153/252d - 9 d = 306/252d - 18 54/252d = 18 d = 84

OpenStudy (dumbcow):

i like your way better lalaly

OpenStudy (lalaly):

Thankyouuuu xD

OpenStudy (anonymous):

Yeah. Lana's is elegant. Just 1 variable.

OpenStudy (anonymous):

alright now i got that answer left off the 5thyear

OpenStudy (anonymous):

Hmm?

OpenStudy (lalaly):

lol i meant :D hehe